OpenStudy (anonymous):

Given P(8,0), Q(-4,9). Write an equation for PQ

OpenStudy (anonymous):

PLEASE HELP ME SOLVE IT!!

OpenStudy (anonymous):

first you want to find the slope: \[m=\frac{ y_{2}-y _{1} }{ x _{2}-x_{1} }\] once you have the slope you need to plug the numbers into this: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (amistre64):

you would also have to restrict the domain if its spose to be only that line segment ...

OpenStudy (amistre64):

if this is with vectors, then you would define the slope as a vector V, and anchor it to a point P such that: x = Px + t Vx y = Py + t Vy and t = 0 to 1 is generally how that would play out